OpenStudy (anonymous):

enlighten me: If f(x) = (x^3)+x and h(x) is the inverse of f(x), then h'(2) is ??

OpenStudy (anonymous):

The inverse of f(x) = h(x) = 1/((x^3)+x) h'(x) = -3(x^-4) -(x^2) h'(2) = -3(32) - (4) h'(2) = -100

OpenStudy (anonymous):

well...the five answer choices are as follows: a) 1/13 b) 1/4 c) 1 d) 4 e) 13

OpenStudy (anonymous):

I think what is meant here is function inverse h(f(x)) = x, not the multiplicative inverse that you meant.

OpenStudy (anonymous):

the question was stated as it appears on the worksheet, its basically asking for the derivative of the inverse of the function f(x)

OpenStudy (anonymous):

the answer is 1/6 i guess, follow my next post

OpenStudy (anonymous):

i thought it was 1/13...if the answer choices mean anything at all then 1/6 is incorrect

OpenStudy (anonymous):

this is what you do. the formula for finding the derivative of an inverse function is \[(f^{-1}(x))'=\frac{1}{f'(f^{-1}(x))}\]\]

OpenStudy (anonymous):

so first find \[f^{-1}(2)\] by solving \[x^3+x=2\] \[x=1\]

OpenStudy (anonymous):

then find \[f'(x)=3x^2+1\] then find \[f'(1)=4\] and so your answer is \[(f^{-1}(2))'=\frac{1}{4}\]

OpenStudy (anonymous):

hope steps are clear

OpenStudy (anonymous):

ok, I couldn't find any better solution. write \[f(x) = y, x = h(y);\] then the equation becomes: \[y = h^3(y) + h(y)\] h is differentiable (given), hence differentiate to get: \[1 = (3 h^2(y) + 1 ) h'(y)\] plug y = 2 \[1 = (3 h^2(2)+1) h'(2)\] we also know that \[h(f(x)) = x \Rightarrow h(x^3+x) = x \Rightarrow h(1+1) = 1\] now put this in the previous expression and get \[h'(2) = 1/(3\times1 + 1) = \frac{1}{4}\] my last calculation was wrong, sorry